OpenStudy (anonymous):

find the values of x and y 0.50x+0.20y=3.90

OpenStudy (lgbasallote):

if you want the numerical value of each then you need another equation

OpenStudy (lgbasallote):

otherwise it's impossible..there are infinite possibilities

OpenStudy (anonymous):

oh it also has the equation x+y=9.0

OpenStudy (lgbasallote):

now it's possible

OpenStudy (lgbasallote):

0.5x + 0.2y = 3.9 x + y = 9.0 you wouldnt mind if i multiply the first equation by 10 to make it look simple would you 5x + 2y = 39 x + y = 9

OpenStudy (lgbasallote):

do you know how to solve that system now?

OpenStudy (anonymous):

then you must use the substitution rule and now you will have y=9 -x and now sub this y into your first equation to get 0.5x +0.2(9-x) = 3.9 and now we will just solve for x and obtain 0.5x + 1.8 -0.2x=3.9 --> 0.3x =2.1 and x=2.1/0.3 = 7 now x=7 and now let us plug this into our third equation y = 9- x and we know x=7 so we we get y=2 therefore our solution is (7,2) :)